OpenStudy (anonymous):

Two numbers, x and y, are randomly chosen from the interval [0,1]. What is the expected value of (x+y)?

OpenStudy (mathmate):

Hint: Find E(x) and E(y) separately. Then use E(X+Y)=E(X)+E(Y)

OpenStudy (anonymous):

Well, I don't really understand what it means by the interval 0,1

OpenStudy (mathmate):

Interval [0,1] is a set of all possible real values between 0 and 1, inclusively. Since it says randomely chosen from the interval, so the random variable is a uniform distribution between 0 and 1. The probability distribution looks like the following:  The function of the distribution is therefore \( P(x)=1\ if\ 0 \le x \le 1\ and \) \(\ \ \ \ \ \ = 0\ elsewhere \) The expected value of a probability distribution p(x) is \( \int_{-\infty}^{\infty}xP(x)dx \\ \) which for a uniform distribution [a,b] is simply the mean of a and b. Hope this gets you started.

OpenStudy (anonymous):

Yeah, I already got the question sorted out ;)

OpenStudy (anonymous):

Thanks anyways though

OpenStudy (mathmate):

yw!